At the beginning of its current fiscal year, Willie Corps balance sheet showed assets of $12,500 and liabilities of $6,000.
Answer #1

Beginning assets = $12,500

Beginning liabilities = $6,000

Beginning stockholders' equity = Beginning assets - Beginning liabilities

= 12,500 - 6,000

= $6,500

Ending liabilities = Beginning liabilities - Decrease in liabilities

= 6,000 - 1,000

= $5,000

Ending assets = $7,000

Ending stockholders' equity = Ending assets - Ending liabilities

= 7,000 - 5,000

= $2,000

Ending stockholders' equity = Beginning stockholders' equity + Net income - Dividend

2,000 = 6,500 + 2,600 - Dividend

Dividend = $7,100

Assets = Liabilities + PIC + RE
Beginning 12,500 = 6,000 + 2,000 + 4,500
Changes -5,500 = -1,000 + + 2,600
= + + -7,100
Ending 7,000 = 5,000 + 2,000 + 0

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.